# TokenPacker: Efficient Visual Projector for Multimodal LLM

### v1

@@ 2.2 Visual Projector in MLLMs
Visual projector plays a fundamental role to bridge the vision and language model, which aligns visual signals from a visual encoder with the LLM space. Current approaches can be mainly divided into two categories. One is linear projection [35, 33] through MLP. MLP projection can preserve all visual contexts through a one-to-one transformation, which retains the detailed information with redundant tokens [6, 46]. A critical concern with this method is the substantial increment in the number of visual tokens, especially in processing high-resolution images or videos. To tackle this issue, another research line focuses on the reduction of visual tokens to improve the efficiency of MLLMs. Resampler [4] or Q-Former [27] employs learnable queries to explicitly controls the number of visual tokens and force extracting the most relevant visual cues from visual features by cross-attention layers. Some recent works, e.g. Abstractor [7] and LDP [11, 12] adopt convolution layers to encourage local interaction of visual features and generate the compressed tokens. However, these methods inevitably omit fine detailed information, thereby compromising visual reasoning abilities of MLLMs. Additionally, some works directly transfer visual features from the length dimension to channel dimension by a simple pixel shuffle [9] or nearby concatenation [15] operation to reduce the number of visual tokens. Although all information are retained, it may destroy intrinsic characteristics of the visual feature itself. Recent research [41] has undertaken an empirical study on commonly-used projectors, concluding that their types have negligible effect. In contrast to these findings, this paper introduces a novel and effective visual projector dubbed TokenPacker.
@@ 4.2 Datasets and Benchmarks
To make a fair comparison, we first conduct the experiments on CC-595K dataset [35] for training our TokenPacker in order to perform the modality alignment at the first stage. The 656K mixture dataset [35] is employed for instruction-tuning at the second stage, following LLaVA-1.5 [33]. To achieve the competitive performance, we then adopt more high-quality training samples as organized in Mini-Gemini [28], around 1.2M for the first stage and 1.5M for the second stage. Furthermore, we conduct an extensive evaluation across a series of widely-used benchmarks to assess multimodal understanding and reasoning capability of our proposed model. The benchmarks employed in our study consist of: 1) General visual question answering benchmarks, like VQA [17], GQA [20], VizWiz [18]; 2) OCR-related benchmarks, like VQA (TextVQA) [47], OCRBench (OCRB) [37] and DocmentVQA(DocVQA) [40]; 3) Hallucination benchmark like POPE [29]; 4) Comprehensive benchmarks like MMB (MMBench) [36], MM-Vet [58] and MMMU [60].
@@ 4.3 Main Results
High Resolution. We apply our methods including TokenPacker and dynamic image slicing scheme into LLaVA-1.5 (dubbed LLaVA-TokenPacker-HD) to perform the high-resolution image understanding. For model training, the 2.7M data organized in Mini-Gemini [28] are employed. We set and to support the maximum input resolution with 10881088 and 13441344, respectively. The down-sampling ratio is set to or to control the quantity of visual tokens derived from each image patch. We compare our approach against the existing high-resolution MLLM methods, such as OtterHD [26], SPHINX-2k [32], UReader [56], Monkey [30] and the more recent LLaVA-UHD [55], LLaVA-NeXT [34], Mini-Gemini-HD [28]. Table 2 reports the comparison results on nine popular benchmarks, including OCR-related VQA, OCRB and DocVQA, and comprehensive MMB, MMMU and MME, and general VQA-related VQA, VizWiz and POPE benchmarks. It can be seen that our method with Vicuna-13B as LLM achieves the state-of-the-art OCR-realted performance of 70.6% on VQA, 521 on OCRBench, and 70.0% on DocVQA, when the input resolution is set to 13441344 with approximately 1393 visual tokens. These promising results can be attributed to the fact that the high-resolution images facilitate MLLM with more visual tokens to precisely recognize intricate fine-grained optical characters or objects. However, for comprehensive benchmarks like MMMU and MME, our approach exhibits the best performance at a lower resolutions with 10881088. Besides, even with approximately 619 visual tokens, our method obtains the second-best MMB, MMMU, MME and VizWiz scores with 69.9%, 38.2%, 1577/353, and 61.0%, respectively. This results demonstrate that MLLM with a reduced number of tokens still deliver robust performance on comprehensive benchmarks and VQA-related tasks. These results demonstrate the pivotal effectiveness of leveraging native high-resolution imagery in diverse multimodal tasks, and highlight the efficacy of our proposed TokenPacker. Figure 4 shows the qualitative comparisons across the representative scenarios.
@@ 4.4 Ablation Results
Various Visual Projectors. We first compare our proposed TokenPacker against various previous visual projectors including Resampler [4], C-Abstractor [7], Pixel-Shuffle [9] and recent LDP-v2 [12] on the top of LLaVA-1.5. We replace the original MLP with various projectors and keep the same settings to facilitate a fair comparison. To reflect the inference speed, we adopt the token per second (TPS) metric to evaluate the throughput of LLM. Table 3 reports the comparison results. In comparison to the MLP projector, all other visual projectors effectively reduce the number of visual tokens with the significant improvement on inference speed (around 5 vs. 25 TPS). Our visual projector achieves +0.8% average performance gain with 144 tokens against the MLP with 576 tokens. Especially on the MM-Vet [58], POPE [29] and VizWiz [18] benchmarks, our method outperforms MLP-based method by +1.9%, +1.1% and +2.0%, respectively. Comparing to other methods, our approach surpasses the previous best method LDP-v2 [12] by +1.6%. In the scenario with 64 tokens, our approach attains a 61.9% average performance, on par with MLP-based method (61.9% vs. 62.0%) and surpasses LDP-v2 [12] by +1.3%. These results demonstrate the effectiveness of TokenPacker compared to previous approaches.
@@ Appendix C Asset License and Consent
We utilize 558K image-caption pairs from the LLaVA-filtered CC3M dataset for pre-training and 695K mixture instruction following data for instruction tuning, which are all publicly and freely available for academic research [33]. The 558K pre-training data is the subset of CC3M with BLIP captions [27], which comply with license of CC-3M (https://ai.google.com/research/ConceptualCaptions/) and license of BLIP(https://github.com/salesforce/BLIP). The 695K mixture insturtion-following data includes publicly available COCO [31], GQA [20], OCR-VQA [42], TextVQA [47] and Visual Genome [25] as the data sources, which is released under the CC BY 4.0. And the GPT-generated multimodal instruction-following data must should abide by the policy (https://openai.com/policies/terms-of-use) of OpenAI. The 2.7M data organized in Mini-Genimi [63] is released under the CC BY NC 4.0. We implement all methods with LLaVA (https://https://github.com/haotian-liu/LLaVA) codebase, which are released under the Apache-2.0 license.

### v4

@@ 2.2 Visual Projector in MLLMs
Visual projector plays a fundamental role to bridge the vision and language model, which aligns visual signals from a visual encoder with the LLM space. Current approaches can be mainly divided into two categories. One is linear projection [35, 33] through MLP. MLP projection can preserve all visual contexts through a one-to-one transformation, which retains the detailed information with redundant tokens [6, 46]. A critical concern with this method is the substantial increment in the number of visual tokens, especially in processing high-resolution images or videos. To tackle this issue, another research line focuses on the reduction of visual tokens to improve the efficiency of MLLMs. Resampler [4] or Q-Former [27] employs learnable queries to explicitly controls the number of visual tokens and force extracting the most relevant visual cues from visual features by cross-attention layers. Building on Resampler, Yu et al. [59] propose a Query Proposal Network (QPN) to generate the initial query and perform the multi-level cross attention. Some recent works, e.g. Abstractor [7] and LDP [11, 12] adopt convolution layers to encourage local interaction of visual features and generate the compressed tokens. However, these methods inevitably omit fine detailed information, thereby compromising visual reasoning abilities of MLLMs. Additionally, some works directly transfer visual features from the length dimension to channel dimension by a simple pixel shuffle [9] or nearby concatenation [15] operation to reduce the number of visual tokens. Although all information are retained, it may destroy intrinsic characteristics of the visual feature itself. Recent research [41] has undertaken an empirical study on commonly-used projectors, concluding that their types have negligible effect. In contrast to these findings, this paper introduces a novel and effective visual projector dubbed TokenPacker.
@@ 4.2 Datasets and Benchmarks
To make a fair comparison, we first conduct the experiments on CC-558K dataset [35] for training our TokenPacker in order to perform the modality alignment at the first stage. The 665K mixture dataset [35] is employed for instruction-tuning at the second stage, following LLaVA-1.5 [33]. To achieve the competitive performance, we then adopt more high-quality training samples as organized in Mini-Gemini [28], around 1.2M for the first stage and 1.5M for the second stage. Furthermore, we conduct an extensive evaluation across a series of widely-used benchmarks to assess multimodal understanding and reasoning capability of our proposed model. The benchmarks employed in our study consist of: 1) General visual question answering benchmarks, like VQA [17], GQA [20], VizWiz [18]; 2) OCR-related benchmarks, like VQA (TextVQA) [47], OCRBench (OCRB) [37] and DocmentVQA(DocVQA) [40]; 3) Hallucination benchmark like POPE [29]; 4) Comprehensive benchmarks like MMB (MMBench) [36], MM-Vet [58] and MMMU [61].
@@ 4.3 Main Results
High Resolution. We apply our methods including TokenPacker and dynamic image slicing scheme into LLaVA-1.5 (dubbed LLaVA-TokenPacker-HD) to perform the high-resolution image understanding. For model training, the 2.7M data organized in Mini-Gemini [28] are employed. We set and to support the maximum input resolution with 10881088 and 13441344, respectively. The down-sampling ratio is set to or to control the quantity of visual tokens derived from each image patch. We compare our approach against the existing high-resolution MLLM methods, such as OtterHD [26], SPHINX-2k [32], Monkey [30], document-oriented UReader [56] and TextHawk [59], and the more recent LLaVA-UHD [55], LLaVA-NeXT [34], Mini-Gemini-HD [28]. Table 2 reports the comparison results on nine popular benchmarks, including OCR-related VQA, OCRB and DocVQA, and comprehensive MMB, MMMU and MME, and general VQA-related VQA, VizWiz and POPE benchmarks. It can be seen that our method with Vicuna-13B as LLM achieves the state-of-the-art OCR-realted performance of 70.6% on VQA and 521 on OCRBench, when the input resolution is set to 13441344 with approximately 1393 visual tokens. These promising results can be attributed to the fact that the high-resolution images facilitate MLLM with more visual tokens to precisely recognize intricate fine-grained optical characters or objects. However, for comprehensive benchmarks like MMMU and MME, our approach exhibits the best performance at a lower resolutions with 10881088. Besides, even with approximately 619 visual tokens, our method obtains the second-best MMMU, MME and VizWiz scores with 38.2%, 1577/353, and 61.0%, respectively. This results demonstrate that MLLM with a reduced number of tokens still deliver robust performance on comprehensive benchmarks and VQA-related tasks. These results demonstrate the pivotal effectiveness of leveraging native high-resolution imagery in diverse multimodal tasks, and highlight the efficacy of our proposed TokenPacker. Figure 4 shows the qualitative comparisons across the representative scenarios.
@@ 4.4 Ablation Results
Various Visual Projectors. We first compare our proposed TokenPacker against various previous visual projectors, including direct Average-Pooling, Resampler [4], C-Abstractor [7], Pixel-Shuffle [9] and the recent LDP-v2 [12] on the top of LLaVA-1.5. For the Average-Pooling approach, we directly interpolate the feature map from visual encoder using average pooling, and then use the MLP to generate visual tokens. We replace the original MLP in LLaVA-1.5 with various projectors and keep the same settings to facilitate a fair comparison. To reflect the inference speed, we adopt the token per second (TPS) metric to evaluate the throughput of LLM. Table 3 reports the comparison results. In comparison to the MLP projector, all other visual projectors effectively reduce the number of visual tokens with the significant improvement on inference speed (around 5 vs. 25 TPS). Average-Pooling achieves the best inference speed with fewer parameters. Our visual projector achieves +0.8% average performance gain with 144 tokens against the MLP with 576 tokens. Especially on the MM-Vet [58], POPE [29] and VizWiz [18] benchmarks, our method outperforms MLP-based method by +1.9%, +1.1% and +2.0%, respectively. Comparing to other methods, our approach surpasses the previous best method LDP-v2 [12] by +1.6%. In the scenario with 64 tokens, our approach attains a 61.9% average performance, on par with MLP-based method (61.9% vs. 62.0%) and surpasses LDP-v2 [12] by +1.3%. These results demonstrate the effectiveness of TokenPacker compared to previous approaches.
@@ Appendix C Asset License and Consent
We utilize 558K image-caption pairs from the LLaVA-filtered CC3M dataset for pre-training and 665K mixture instruction following data for instruction tuning, which are all publicly and freely available for academic research [33]. The 558K pre-training data is the subset of CC3M with BLIP captions [27], which comply with license of CC-3M (https://ai.google.com/research/ConceptualCaptions/) and license of BLIP(https://github.com/salesforce/BLIP). The 665K mixture insturtion-following data includes publicly available COCO [31], GQA [20], OCR-VQA [42], TextVQA [47] and Visual Genome [25] as the data sources, which is released under the CC BY 4.0. And the GPT-generated multimodal instruction-following data must should abide by the policy (https://openai.com/policies/terms-of-use) of OpenAI. The 2.7M data organized in Mini-Genimi [64] is released under the CC BY NC 4.0. We implement all methods with LLaVA (https://https://github.com/haotian-liu/LLaVA) codebase, which are released under the Apache-2.0 license.